# Alternating Minimization Schemes for Computing Rate-Distortion-Perception Functions with $f$-Divergence Perception Constraints

### v1

@@ Appendix F Proof of Theorem 5
Furthermore, let $\Lambda_{s,v[q_{Y}]}$ be the set defined as
$\Lambda_{s,v[q_{Y}]}\triangleq\big{\{}\lambda\in\mathbb{R}^{|\mathcal{X}|}:\forall x\in\mathcal{X},\lambda(x)\geq 0\land\forall y\in\mathcal{X},~{}\sum_{x\in\mathcal{X}}p_{X}(x)\lambda(x)A(x,,s)\leq 1\big{\}}$
@@
$\begin{split}R(D,P)\geq&-\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\frac{1}{\lambda(x)}\right)-s_{D}D-s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)\tilde{g}(y).\end{split}$
@@
$\begin{split}&I\left(p_{X},Q_{Y|X}\right)+\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\frac{1}{\lambda(x)}\right)+s_{D}D+s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)\tilde{g}(y)\\ &\geq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\log\left(\frac{Q_{Y|X}}{q_{Y}\lambda(x)A(x,,s)}\right)\\ &\geq 1-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\left(\frac{q_{Y}\lambda{x}A(x,,s)}{Q_{Y|X}}\right)\\ &=1-\sum_{y\in\mathcal{X}}q_{Y}\sum_{x\in\mathcal{X}}p_{X}\lambda(x)A(x,,s)\\ &\geq 1-\sum_{y\in\mathcal{X}}q_{Y}=0.\end{split}$
@@
$\begin{split}\lambda(x)=\frac{1}{\sum_{y\in\mathcal{X}}q_{Y}A(x,,s)}\\ \end{split}$
@@
$\begin{split}R(D,P)&\leq\quad I\left(p_{X},\hat{Q}_{Y|X}\right)\\ &=\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)\hat{Q}_{Y|X}(y|x)\log\left(\frac{\hat{Q}_{Y|X}(y,x)}{\sum_{x\in\mathcal{X}}p_{X}(x)\hat{Q}_{Y|X}(y|x)}\right)\\ &=-s_{D}D-s_{P}P+\sum_{y\in\mathcal{X}}\hat{h}c()\log(c())+s_{P}\sum_{y\in\mathcal{X}}q_{Y}(y)\frac{p_{X}(y)}{v[q_{Y}](y)}\partial f\left(\frac{p_{X}(y)}{v[q_{Y}](y)}\right)\\ &\quad-s_{P}\Bigg{\{}\sum_{y\in\mathcal{X}}q_{Y}(y)f\left(\frac{p_{X}(y)}{v[q_{Y}](y)}\right)-P\Bigg{\}}-s_{P}\sum_{j\in\mathcal{X}}\hat{h}(j)A(x,j,s)\\ &=-s_{D}D-s_{P}P+\sum_{y\in\mathcal{X}}\hat{h}c()\log(c())+\hat{W}[\hat{h}].\end{split}$
@@
$\lambda(x)=\left(c_{\max}\sum_{y\in\mathcal{X}}\hat{h}(y)A(x,,s)\right)$
@@ Appendix G Proof of Theorem 6
$\begin{split}\frac{\partial S[h](i)}{\partial h(j)}&=h(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{\partial}{\partial h(j)}\left(\frac{A[S[h]](x,i,s)}{\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s)}\right)+c[h,S[h]](i)\delta_{i,j}\end{split}$
@@
$\begin{split}\frac{\partial(\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s))}{\partial h(j)}\ =\sum_{k\in\mathcal{X}}h(k)\frac{\partial A[S[h]](x,k,s)}{\partial h(j)}+A[S[h]](x,j,s)\end{split}$
$\begin{split}\frac{\partial A(x,i,s)}{\partial h(j)}=-s_{P}A[S[h]](x,i,s)\frac{p_{X}(i)^{2}}{(S[h(i)])^{3}}\partial f\left(\frac{p_{X}(i)}{S[h(i)]}\right)\frac{\partial S[h](i)}{\partial h(j)}\end{split}.$
@@
and defining the matrices $M$ and $\Gamma$ as in (LABEL:matrix:_Newton_M) and (14), we can rewrite the entries of $J[h^{*}]$ as
$\begin{split}J[h^{*}](i,j)&=c[h^{*},h^{*}](i)\left(\delta_{i,j}-\Gamma[h^{*},h^{*}](i)J[h^{*}](i,j)\right)\\ &\quad+\sum_{k\in\mathcal{X}}\Gamma[h^{*},h^{*}](k)M[h^{*},h^{*}](i,k)J[h^{*}](k,j)-M[h^{*},h^{*}](i,j)\\ \end{split}$ (31)
@@
$\begin{split}J[h^{*}]&=C[h^{*},h^{*}](I-\Gamma[h^{*},h^{*}]J[h^{*}])-M[h^{*},h^{*}]+M[h^{*},h^{*}]\Gamma[h^{*},h^{*}]J[h^{*}]\\ &=(C[h^{*},h^{*}]-M[h^{*},h^{*}])(I-\Gamma J[h^{*}])\end{split}$
@@ Appendix I Proof of Lemma 6
$\begin{split}\sum_{i\in\mathcal{X}}M^{*}(i,j)&=\sum_{i\in\mathcal{X}}h^{*}(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[h^{*}](x,i,s)A[h^{*}](x,j,s){}}{(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s))^{2}}\\ &=\sum_{x\in\mathcal{X}}\frac{A[h^{*}](x,j,s)}{\left(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s)\right)^{2}}\left(\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)\right)\\ &=c[h^{*},h^{*}](j)=1.\end{split}$
@@
$1\displaystyle 0<\operatorname{eig}(M^{*})\leq 1\implies 0\leq\operatorname{eig}(I-M^{*})<$ 1
@@
$1\leq\operatorname{eig}(I+(I-M^{*})\Gamma)\implies 0<\operatorname{eig}((I+(I-M^{*})\Gamma^{*})^{-1})\leq 1$
@@
Define $\theta_{\sup}\triangleq\operatorname{eig}_{\max}(I-M^{*})\cdot\operatorname{eig}_{\max}((I+(I-M^{*})\Gamma^{*})^{-1})$. Then, we can show that $10\leq\operatorname{eig}(J[h^{*}])\leq\theta_{\sup}<$ 1 is always verified. The second part of the theorem follows directly from [42, Theorem 5] hence we omit it. This completes the proof.
@@
In order to guarantee $10\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<$ 1, it is sufficient to have
@@
Since in the non-degenerate case we have $10<\theta_{\max}<$ 1, we can construct $\epsilon^{\prime}=\min\left\{s_{P,\max},\frac{1}{f^{\prime\prime}(0)}\left(\frac{1}{\theta_{\max}}-1\right)\right\}$ > 0 and define the set $I_{s_{P}}^{\epsilon^{\prime}}$ where the exponential convergence of the approximate algorithm is guaranteed. This concludes the proof.
@@
We remind that $f=|x-1|$ is the function associated with the TV distance $TV(\cdot||\cdot)$. The first statement can be proved by first establishing the uniform convergence of $f_{n}\to f$ as $n\to\infty$
$\sup_{x\in\mathbb{R}}\left|f_{n}(x)-f(x)\right|$ $=\sup_{x\in\mathbb{R}}\left|\frac{2}{\pi}(x-1)\arctan(n(x-1))-|x-1|\right|$
@@
$=\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{\pi}x\arctan\left(\frac{1}{nx}\right)$
$\leq\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{n\pi}=\frac{2}{n\pi}$
@@
$\lim_{n\to\infty}|D_{f_{n}}(p||q)-TV(p||q)|$ $=\lim_{n\to\infty}\left|\sum_{x\in\mathcal{X}}q(x)\left(f_{n}\left(\frac{p(x)}{q(x)}\right)-f\left(\frac{p(x)}{q(x)}\right)\right)\right|$

### v2

@@ Appendix F Proof of Theorem 5
where $q_{Y}=\sum_{x\in\mathcal{X}}p_{X}Q_{Y|X}$. Furthermore, let $\Lambda_{s,v[q_{Y}]}$ be the set defined as
$\Lambda_{s,v[q_{Y}]}\triangleq\Big{\{}\lambda\in\mathbb{R}^{|\mathcal{X}|}$ $\forall x\in\mathcal{X},\lambda(x)\geq 0\land\forall y\in\mathcal{X},$
$\sum_{x\in\mathcal{X}}p_{X}(x)\lambda(x)A[v[q_{Y}]](x,y,s)\leq 1\Big{\}}.$
@@
$\begin{split}R(D,P)\geq&\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\lambda(x)\right)-s_{D}D\\ &-s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)g(p_{X}(y),v[q_{Y}](y)).\end{split}$
@@
$\begin{split}&I\left(p_{X},Q_{Y|X}\right)+\sum_{x\in\mathcal{X}}p_{X}(x)\log\left(\frac{1}{\lambda(x)}\right)+s_{D}D\\ &\quad+s_{P}\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}(x)Q_{Y|X}(x,y)g(p_{X}(y),v[q_{Y}](y))\\ &\geq\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\log\left(\frac{Q_{Y|X}}{q_{Y}\lambda(x)A[v[q_{Y}]](x,y,s)}\right)\\ &\geq 1-\sum_{(x,y)\in\mathcal{X}^{2}}p_{X}Q_{Y|X}\left(\frac{q_{Y}\lambda(x)A[v[q_{Y}]](x,y,s)}{Q_{Y|X}}\right)\\ &=1-\sum_{y\in\mathcal{X}}q_{Y}\sum_{x\in\mathcal{X}}p_{X}\lambda(x)A[v[q_{Y}]](x,y,s)\\ &\geq 1-\sum_{y\in\mathcal{X}}q_{Y}=0.\end{split}$
@@
[…]
@@
[…]
@@
$\lambda(x)=\left(c_{\max}\sum_{y\in\mathcal{X}}\hat{h}(y)A[v[\hat{h}]](x,y,s)\right)^{-1}$
@@ Appendix G Proof of Theorem 6
$\begin{split}&\frac{\partial S[h](i)}{\partial h(j)}=c[h,S[h]](i)\delta_{i,j}\\ &\quad+h(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{\partial}{\partial h(j)}\left(\frac{A[S[h]](x,i,s)}{\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s)}\right)\end{split}$
@@
$\begin{split}&\frac{\partial(\sum_{k\in\mathcal{X}}h(k)A[S[h]](x,k,s))}{\partial h(j)}=\\ &\qquad\sum_{k\in\mathcal{X}}h(k)\frac{\partial A[S[h]](x,k,s)}{\partial h(j)}+A[S[h]](x,j,s)\end{split}$
$\begin{split}&\frac{\partial A(x,i,s)}{\partial h(j)}=\\ &\qquad-s_{P}A[S[h]](x,i,s)\frac{p_{X}(i)^{2}}{(S[h(i)])^{3}}\partial f\left(\frac{p_{X}(i)}{S[h(i)]}\right)\frac{\partial S[h](i)}{\partial h(j)}\end{split}.$
@@
and defining the matrices $M$ and $\Gamma$ as in (14) and (15), we can rewrite the entries of $J[h^{*}]$ as
$J[h^{*}](i,j)$ $=c[h^{*},h^{*}](i)\left(\delta_{i,j}-\Gamma[h^{*},h^{*}](i)J[h^{*}](i,j)\right)$
$\quad+\sum_{k\in\mathcal{X}}\Gamma[h^{*},h^{*}](k)M[h^{*},h^{*}](i,k)J[h^{*}](k,j)$
$\quad-M[h^{*},h^{*}](i,j)$ (30)
@@
$\begin{split}J[h^{*}]&=C[h^{*},h^{*}](I-\Gamma[h^{*},h^{*}]J[h^{*}])\\ &\quad-M[h^{*},h^{*}]+M[h^{*},h^{*}]\Gamma[h^{*},h^{*}]J[h^{*}]\\ &=(C[h^{*},h^{*}]-M[h^{*},h^{*}])(I-\Gamma J[h^{*}])\end{split}$
@@ Appendix I Proof of Lemma 6
$\sum_{i\in\mathcal{X}}M^{*}(i,j)$
$=\sum_{i\in\mathcal{X}}h^{*}(i)\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[h^{*}](x,i,s)A[h^{*}](x,j,s){}}{(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s))^{2}}$
$=\sum_{x\in\mathcal{X}}p_{X}(x)\frac{A[h^{*}](x,j,s)\left(\sum_{i\in\mathcal{X}}h^{*}(i)A[h^{*}](x,i,s)\right)}{\left(\sum_{k\in\mathcal{X}}h^{*}(k)A[h^{*}](x,k,s)\right)^{2}}$
$=c[h^{*},h^{*}](j)=1.$
@@
$0<\operatorname{eig}(M^{*})\leq 1$ $\implies 0\leq\operatorname{eig}(I-M^{*})<1$
@@
$1\leq\operatorname{eig}(I+(I-M^{*})\Gamma)$ $\implies$
$0<\operatorname{eig}((I+(I-M^{*})\Gamma^{*})^{-1})\leq 1$
@@
Define $\theta_{\sup}\triangleq\operatorname{eig}_{\max}(I-M^{*})\cdot\operatorname{eig}_{\max}((I+(I-M^{*})\Gamma^{*})^{-1})$. Then, we can show that $0\leq\operatorname{eig}(J[h^{*}])\leq\theta_{\sup}<1$ is always verified. The second part of the theorem follows directly from [44, Theorem 5] hence we omit it. This completes the proof.
@@
In order to guarantee $0\leq\{\theta_{a,i}\}_{i\in\mathcal{X}}<1$, it is sufficient to have
@@
Since in the non-degenerate case we have $0<\theta_{\max}<1$, we can construct $\epsilon^{\prime}=\min\left\{s_{P,\max},\frac{1}{f^{\prime\prime}(0)}\left(\frac{1}{\theta_{\max}}-1\right)\right\}$ > 0 and define the set $I_{s_{P}}^{\epsilon^{\prime}}$ where the exponential convergence of the approximate algorithm is guaranteed. This concludes the proof.
@@
We remind that f=|x1|f=|x-1| is the function associated with the TV distance $TV(\cdot||\cdot)$. The first statement can be proved by first establishing the uniform convergence of $f_{n}\to f$ as $n\to\infty$
$\sup_{x\in\mathbb{R}}\left|f_{n}(x)-f(x)\right|$
$=\sup_{x\in\mathbb{R}}\left|\frac{2}{\pi}(x-1)\arctan(n(x-1))-|x-1|\right|$
@@
$=\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{\pi}x\arctan\left(\frac{1}{nx}\right)\leq\sup_{x\in\mathbb{R}^{+}_{0}}\frac{2}{n\pi}=\frac{2}{n\pi}$
@@
$\lim_{n\to\infty}|D_{f_{n}}(p||q)-TV(p||q)|$
$=\lim_{n\to\infty}\left|\sum_{x\in\mathcal{X}}q(x)\left(f_{n}\left(\frac{p(x)}{q(x)}\right)-f\left(\frac{p(x)}{q(x)}\right)\right)\right|$